OpenStudy (hiyathere_20):

Find the solution to x2 = 121.

OpenStudy (benji):

Is the x raised to the power of 2?

OpenStudy (hiyathere_20):

yes

OpenStudy (benji):

Alright, so how do you "undo" a square?

OpenStudy (skullpatrol):

$$\huge x^2=121$$ means what number(s) multiplied by themselves equal 121?

OpenStudy (benji):

If \[2^{2}=4 \]what does \[\sqrt{4}=?\]

OpenStudy (hiyathere_20):

16?

OpenStudy (hiyathere_20):

wait, 2?

OpenStudy (benji):

Yes, 2, so if we apply that method to the 121, what do you get?

OpenStudy (skullpatrol):

$$\huge x • x = 121$$

OpenStudy (skullpatrol):

10 • 10 = ? 12 • 12 = ?

OpenStudy (hiyathere_20):

OHHH SO It's 11?

OpenStudy (skullpatrol):

Yes, that is one of them. What is the other?

OpenStudy (hiyathere_20):

There should only be answer. And mines is positive 11

OpenStudy (hiyathere_20):

a. ±121 b. -11 c. ±11 d. 11

OpenStudy (skullpatrol):

What about –11 times –11?

OpenStudy (hiyathere_20):

-22?

OpenStudy (skullpatrol):

That's plus.

OpenStudy (skullpatrol):

x^2 means x times x.

OpenStudy (hiyathere_20):

okay so what is the answer 11 ? since that equals 121...

OpenStudy (skullpatrol):

There are two answers that will make $$\huge x^2 = 121$$ a true statement.

OpenStudy (hiyathere_20):

Omg but there needs to be only one answer for this question........

OpenStudy (skullpatrol):

11(11) = 121 -11(-11) = 121

OpenStudy (skullpatrol):

$$\huge x= \pm11$$

OpenStudy (skullpatrol):

There is only one answer to this question. That answer has two numbers: +11 and –11